# Curvature Tuning: Provable Training-free Model Steering From a Single Parameter

### v4

@@ Abstract
The scaling of model and data sizes has reshaped the AI landscape, establishing finetuning pretrained models as the standard paradigm for solving downstream tasks. However, dominant finetuning methods typically rely on weight adaptation, often lack interpretability, and depend on heuristically chosen hyperparameters. In this paper, we take a different perspective and shift the focus from weights to activation functions, viewing them through the lens of spline operators. We propose Curvature Tuning (CT), an interpretable and principled steering method that modulates a model’s decision boundary by injecting a single hyperparameter into its activation functions. We show that CT provably adjusts model decision boundary curvature and, more fundamentally, projects a model onto a space of smooth functions—thereby complementing current finetuning methods, whose effect lies primarily in feature adaptation. Making this hyperparameter trainable gives rise to a novel and highly parameter-efficient finetuning method. Empirically, CT improves both generalization and robustness. For example, it boosts downstream accuracy of ResNet-50/152 by 7.14%/8.46% over linear probing and 4.64%/1.70% over LoRA across 12 datasets, and improves robust accuracy on the $\ell_{\infty}$ benchmark from RobustBench by 1032.64%/1494.46%. Our code is available at https://github.com/Leon-Leyang/curvature-tuning.
@@ 1 Introduction
Full finetuning, the process of steering [^1] In this paper, we use steering as a general term for tuning a model, including both training-based and non-training-based methods. We use finetuning to refer specifically to steering methods that adapt the model’s parameters via training. a pretrained model by adapting all its parameters to downstream datasets, was once the primary approach for transferring knowledge. While it effectively enhances generalization [6] and robustness [7], it is computationally expensive at large model scales. To mitigate this, parameter-efficient finetuning (PEFT) methods such as Serial Adapter [8] and LoRA [9] have been introduced, which finetune only a small subset of parameters. However, these approaches usually lack interpretability and principled design. For instance, they treat the model as a black box, making it unclear how the model is steered for downstream tasks. Consequently they usually rely on heuristic choices—such as LoRA’s rank, placement, and initialization—with minimal theoretical guidance. This leads to a natural question: how can we construct principled steering solutions addressing both efficiency and interpretability?
@@
Grounded in the spline interpretation of deep networks [12, 13], we propose Curvature Tuning (CT), a steering method that provably modulates a model’s decision boundary curvature by injecting a single hyperparameter $\beta$ into the activation function, as shown in Fig. 1. Additionally, allowing $\beta$ to be trained leads to a novel finetuning method. We highlight four key advantages of CT below:
@@
CT is highly parameter-efficient. As a steering method, CT introduces only one (hyper)parameter per network. As a finetuning method, Trainable CT still uses significantly fewer parameters than LoRA with rank one, requiring only 0.58% to 59.05% of the parameters used by LoRA in our experiments.
CT improves both generalization and robustness. For example, Trainable CT boosts transfer accuracy of ResNet-50/152 by 7.14%/8.46% over linear probing and 4.64%/1.70% over LoRA with rank one across 12 downstream datasets. CT improves robust accuracy of ResNet-50/152 by 1032.64%/1494.46% on the $\ell_{\infty}$ benchmark from RobustBench.
In summary, our key contributions are both theoretical and empirical. Theoretically, we propose Curvature Tuning (CT) and show that it provably modulates the model’s decision boundary curvature, by projecting the model onto a space of smooth functions. Empirically, we introduce CT as a steering method and Trainable CT as a finetuning method, demonstrating improved generalization across five models and 12 downstream datasets, as well as robustness gains on the RobustBench benchmark [14].
The remainder of this paper is organized as follows: Section 2 reviews current finetuning techniques and introduces relevant spline concepts, the foundation for our method. Section 3 details our proposed method and its theoretical guarantees. Section 4 presents experimental results, and Section 5 summarizes our findings and potential future directions.
@@ 2 Background
This section first reviews current finetuning techniques and their limitations (Section 2.1), then introduces spline theory and its connections to deep networks as the foundation for CT (Section 2.2).
@@ 2.1 Current finetuning techniques and limitations
Finetuning refers to steering a pretrained model to improve its performance on downstream tasks through training. Initially, the common practice was to continue training all model parameters—a process known as full finetuning. Notable examples include GPT [6] and DINO [15]. However, as model sizes have grown, full finetuning has become increasingly costly and impractical, especially given the limited size of many downstream datasets. Given these challenges, parameter-efficient finetuning (PEFT) methods were developed to mitigate the cost while maintaining effectiveness.
We follow the categorization of Han et al. [16], which groups PEFT methods into four main categories. Additive PEFT introduces additional trainable parameters to the pretrained model, training only these new parameters during finetuning. Examples include Serial Adapter [8], Prefix-tuning [17], (IA)3 [18] and RoAd [19]. Selective PEFT identifies a subset of existing parameters for finetuning, with examples such as U-Diff pruning and S-Diff pruning [20]. Reparameterized PEFT decomposes pretrained weights into low-rank matrices, finetuning only the low-rank components, which are converted back during inference; examples include LoRA [9] and DyLoRA [21]. Hybrid PEFT combines multiple PEFT approaches, such as UniPELT [22] and S4 [23].
While PEFT methods differ in the parameters they update, they all adapt model weights and operate on learned features—an approach that often relies on heuristic tuning. For example, LoRA requires decisions about adapter placement [24], rank [21, 25], scaling [26], and initialization [27]. In contrast, as described in Section 3, CT introduces only a single hyperparameter into the activation functions that provably modulates the decision boundary curvature, offering a more interpretable alternative that instead operates on the model’s underlying function space without changing model weights.
@@ 2.2 The spline formulation of deep networks
In this subsection, we review relevant concepts in splines, which provide a mathematical framework for understanding the relationship between piecewise-affine functions and deep networks (DNs).
@@
The key result underpinning our study is that many DN layers—such as fully connected and convolutional layers, and convex piecewise-affine activations (e.g., ReLU, max pooling, or maxout)—can be exactly represented as max-affine spline functions [13] (further details in Appendix A).
Now that we have reviewed existing finetuning methods and their limitations, and introduced the necessary spline-based foundations, we proceed to present our proposed method in Section 3.
@@ 3 Curvature Tuning (CT): a provable method for model steering
In this section, we introduce our proposed method, Curvature Tuning (CT). We begin by motivating the benefits of modulating decision boundaries as a model steering technique. Then, we dive into CT’s construction in Section 3.1 with implementation details in Section 3.2. Additional theoretical intuition is provided in Section 3.3. Extensive experimental validation is presented in Section 4.
(a) Baseline
@@
Figure 2: Toy example illustrating how modulating model’s activation functions steers decision boundaries curvature. The model is a 2-layer MLP with hidden width 7; (a) baseline trained for 4000 steps, then fine-tuned for another 4000 steps using (b) LoRA ($r=1$, $\alpha=1$) and (c) Trainable CT. CT achieves near-optimal approximation by smoothing the decision boundary of the pretrained model, whereas LoRA only operates on the model parameters, without changing the model’s underlying geometry.
[…]
3.1 The $\beta$-VQ inference framework
This section builds upon the max-affine spline formulation from Eq. 2 to construct a model steering method operating on the model’s activation functions. By inspecting Eq. 2, we observe that the mapping remains affine within each (implicitly defined) region where the pointwise maximum does not change. Specifically, for any input $\mathbf{x}$ where $\arg\max_{r=1\dots R}\big{(}\langle\mathbf{A}_{r,\cdot},\mathbf{x}\rangle+\mathbf{b}_{r}\big{)}$ remains constant, all such inputs belong to the same region, as they share the same affine mapping. The non-linearity of the function arises when transitioning between these regions.
Smoothing the non-linearity by smoothing the spline region assignment process. Instead of going from one affine mapping to another in an abrupt fashion (whenever crossing that hyperplane), one may consider a smoother transition. There are two common practices to achieve that goal.
@@ 3.1 The $\beta$-VQ inference framework
1. VQ Inference Step (region selection): Determine the affine transformation that maximizes the output, which can be viewed as a vector quantization (VQ) process. The decision is encoded in a one-hot selection variable $\mathbf{t}\in\mathbb{R}^{R}$, where $R$ is the number of input region partitions of the max-affine spline function, and the $r^{*}$-th entry is set to 1, where:
@@
Using the computation step in Eq. 4 and a ReLU activation function, switching from $\beta=1$ to $\beta=0.5$ is provably equivalent to replacing ReLU with the Sigmoid Linear Unit (SiLU). In the limit as $\beta\to 0$, the activation function becomes linear—thus making the entire input-output mapping of the network linear as well.
Smoothing the nonlinearity by smoothing the max. As previously mentioned, there is an alternative way to smooth the max-affine spline mapping from Eq. 2. Instead of relying on a soft region assignment, we can instead directly smooth the maximum function. It is already well known that smoothing the maximum operator leads to the log-sum-exp operator (i.e. SoftPlus). Hence, the mapping from Eq. 2 now becomes
@@
where we parameterized the mapping so that its behavior is akin to Eq. 5, a value of $\beta\to 1$ recovers the original affine spline activation, e.g., ReLU.
The crucial observation we make is that both parameterizations tend to shift the mean of the output of the unit either by a negative factor (for Eq. 5) or by a positive factor (for Eq. 7). This means that in very deep models, varying $\beta$ with either parameterization produces a shift in the decision boundary or regression that cannot be recovered unless the parameters are trained once again, which we are trying to avoid. As a result, as detailed in Section 3.2, our implementation combines the two parameterizations in a weighted manner to mitigate this bias, as Appendix Fig. 3 illustrates.
@@ 3.2 Implementation of CT
We begin by presenting the core activation that gives CT its expressive power—referred to as CT Unit (CTU). The activation is obtained by combining the two parameterizations discussed in Section 3.1, in order to mitigate the mean shift introduced by each parameterization individually:
@@
Steering vs Trainable CT. We provide two implementations of CT differing in how CTU is applied. The first, denoted CT, replaces all ReLUs in the network with CTUs using a fixed $c=0.5$ and a shared $\beta\in[0,1]$. This version is highly parameter-efficient—introducing only a single hyperparameter—and does not require backpropagation, making it suitable as a steering method.
The second, referred to as Trainable CT, also replaces all ReLUs with CTUs but assigns each output neuron its own trainable pair $(\beta,c)$, optimized via backpropagation. This version serves as a finetuning method: while it introduces additional parameters, the increase is modest compared to methods like LoRA and it yields better performance, as shown in Section 4.3. Code for both implementations is provided in Appendix D.

### v5

@@ Abstract
The scaling of model and data sizes has reshaped the AI landscape, establishing finetuning pretrained models as the standard paradigm for solving downstream tasks. However, dominant finetuning methods typically rely on weight adaptation, often lack interpretability, and depend on heuristically chosen hyperparameters. In this paper, we take a different perspective and shift the focus from weights to activation functions, viewing them through the lens of spline operators. We propose Curvature Tuning (CT), an interpretable and principled steering method that modulates a model’s decision boundary by injecting a single hyperparameter into its activation functions. We show that CT provably adjusts model decision boundary curvature and, more fundamentally, projects a model onto a space of smooth functions—thereby complementing current finetuning methods, whose effect lies primarily in feature adaptation. Making this hyperparameter trainable gives rise to a novel and highly parameter-efficient finetuning method. Empirically, CT improves both generalization and robustness. For example, it boosts downstream accuracy of ResNet-50/152 by 8.59%/8.34% over linear probing and 4.64%/1.70% over LoRA across 12 datasets, and improves robust accuracy on the $\ell_{\infty}$ benchmark from RobustBench by 1032.64%/1494.46%. Our code is available at https://github.com/Leon-Leyang/curvature-tuning.
@@ 1 Introduction
Full finetuning, the process of steering [^1] In this paper, we use steering as a general term for tuning a model, including both training-based and non-training-based methods. We use finetuning to refer specifically to steering methods that adapt the model’s parameters via training. a pretrained model by adapting all its parameters to downstream datasets, was once the primary approach for transferring knowledge. While it effectively enhances generalization [34] and robustness [21], it is computationally expensive at large model scales. To mitigate this, parameter-efficient finetuning (PEFT) methods such as Serial Adapter [19] and LoRA [20] have been introduced, which finetune only a small subset of parameters. However, these approaches usually lack interpretability and principled design. For instance, they treat the model as a black box, making it unclear how the model is steered for downstream tasks. Consequently, they usually rely on heuristic choices—such as LoRA’s rank, placement, and initialization—with minimal theoretical guidance. This leads to a natural question: how can we construct principled steering solutions addressing both efficiency and interpretability?
@@
Grounded in the spline interpretation of deep networks [30, 2], we propose Curvature Tuning (CT), a steering method (denoted as S-CT) that provably modulates a model’s decision boundary curvature by injecting a single hyperparameter $\beta$ into the activation function, as shown in Fig.˜1. Additionally, allowing $\beta$ to be trained leads to a novel finetuning method (denoted as T-CT). We highlight four key advantages of CT below:
@@
CT is highly parameter-efficient. As a steering method, S-CT introduces only one (hyper)parameter per network. As a finetuning method, T-CT still uses significantly fewer parameters than LoRA with rank one, requiring only 0.58% to 59.09% of the parameters used by LoRA in our experiments.
CT improves both generalization and robustness. For example, T-CT boosts transfer accuracy of ResNet-50/152 by 8.59%/8.34% over linear probing and 4.64%/1.70% over LoRA with rank one across 12 downstream datasets. S-CT improves robust accuracy of ResNet-50/152 by 1032.64%/1494.46% on the $\ell_{\infty}$ benchmark from RobustBench.
In summary, our key contributions are both theoretical and empirical. Theoretically, we propose CT and show that it provably modulates the model’s decision boundary curvature, by projecting the model onto a space of smooth functions. Empirically, we introduce S-CT as a steering method and T-CT as a finetuning method, demonstrating improved generalization across six models and 12 downstream datasets, as well as robustness gains on the RobustBench benchmark [8].
The remainder of this paper is organized as follows: Section˜2 reviews current finetuning techniques and introduces relevant spline concepts, the foundation for our method. Section˜3 details our proposed method and its theoretical guarantees. Section˜4 presents experimental results, and Section˜5 summarizes our findings and potential future directions.
@@ 2 Background
This section first reviews current finetuning techniques and their limitations (Section˜2.1), then introduces spline theory and its connections to deep networks as the foundation for CT (Section˜2.2).
@@ 2.1 Current finetuning techniques and limitations
Finetuning refers to steering a pretrained model to improve its performance on downstream tasks through training. Initially, the common practice was to continue training all model parameters—a process known as full finetuning. Notable examples include GPT [34] and DINO [4]. However, as model sizes have grown, full finetuning has become increasingly costly and often impractical, particularly when downstream datasets are small. Given these challenges, parameter-efficient finetuning (PEFT) methods were developed to mitigate the cost while maintaining effectiveness.
We follow the categorization of Han et al. [15], which groups PEFT methods into four main categories. Additive PEFT introduces additional trainable parameters to the pretrained model, training only these new parameters during finetuning. Examples include Serial Adapter [19], Prefix-tuning [24], (IA)3 [26], and RoAd [25]. Selective PEFT identifies a subset of existing parameters for finetuning, with examples such as U-Diff pruning and S-Diff pruning [14]. Reparameterized PEFT decomposes pretrained weights into low-rank matrices, finetuning only the low-rank components, which are converted back during inference; examples include LoRA [20] and DyLoRA [36]. Hybrid PEFT combines multiple PEFT strategies, such as UniPELT [29] and S4 [5].
While PEFT methods differ in the parameters they update, they all adapt model weights and operate on learned features—an approach that often relies on heuristic tuning. For example, LoRA requires decisions about adapter placement [13], rank [36, 6], scaling [22], and initialization [16]. In contrast, as described in Section˜3, CT introduces only a single hyperparameter into the activation functions that provably modulates the decision boundary curvature, offering a more interpretable alternative that instead operates on the model’s underlying function space without changing model weights.
The problem of learning low-curvature activation functions has been explored in the context of pretraining, where Srinivas et al. [35] exploit a parametric version of Softplus. Our paper differs in two substantial aspects. First, we introduce a more general parametric activation function (CTU) encompassing a wider family of non-linearities, which provides direct applicability to current network architectures and particularly transformers. Second, we are the first to propose modulating activation functions as a novel PEFT direction, allowing to control downstream generalization and robustness without changing the pre-trained model’s learned features.
@@ 2.2 The spline formulation of deep networks
In this subsection, we review relevant concepts in splines, which provide a mathematical framework for understanding the relationship between piecewise-affine functions and deep networks (DN).
@@
The key result underpinning our study is that many DN layers—such as fully connected and convolutional layers, and convex piecewise-affine activations (e.g., ReLU, max pooling, and maxout)—can be exactly represented as max-affine spline operators [^2] Here, operator simply refers to a mapping between vector spaces (i.e., a function $f:\mathbb{R}^{D}\to\mathbb{R}^{K}$) obtained by concatenating $K$ functions like $s$., and that DNs composed of such layers can be represented as affine spline operators [2] (further details in Appendix˜A).
Now that we have reviewed existing finetuning methods and their limitations, and introduced the necessary spline-based foundations, we proceed to present our proposed method in Section˜3.
@@ 3 Curvature Tuning (CT): a provable method for model steering
In this section, we introduce our proposed method, Curvature Tuning (CT). We begin by motivating the benefits of modulating decision boundaries as a model steering technique. Then, we dive into CT’s construction in Section˜3.1 with implementation details in Section˜3.2. Additional theoretical intuition is provided in Section˜3.3. Extensive experimental validation is presented in Section˜4.
(a) Pretrained
@@
Figure 2: Toy example illustrating how modulating model’s activation functions steers decision boundaries curvature. The model is a 2-layer MLP with hidden width 7; (a) baseline pretrained for 4000 steps, then finetuned for another 4000 steps using (b) LoRA ($r=1$, $\alpha=1$) and (c) Trainable CT. CT achieves near-optimal approximation by smoothing the decision boundary of the pretrained model, whereas LoRA only operates on the model parameters, without changing the model’s underlying geometry.
Motivating example.    Consider a toy binary classification problem in $\mathbb{R}^{2}$, whereby the optimal decision boundary separating two classes is given by the unit circle $S^{1}=\{\mathbf{x}\in\mathbb{R}^{2}:\|\mathbf{x}\|_{2}=1\}$, parameterized by the curve $\bm{\gamma}:t\mapsto(\cos 2\pi t,\sin 2\pi t)$, for $t\in{[}0,1{]}$. Let $\sigma(z)=\frac{\exp\left(z\right)}{1+\exp\left(z\right)}$ be the sigmoid. By definition of decision boundary, an optimal network $f:\mathbb{R}^{2}\to\mathbb{R}$ should predict both classes with equal probability $\sigma(f(\bm{\gamma}(t)))=0.5,\forall t\in{[}0,1{]}\iff f(\bm{\gamma}(t))=0,\forall t$. Focusing on the decision boundary, the approximation error $e$ is given by the line integral $e=\int_{\bm{\gamma}}|f(\mathbf{x})|d\mathbf{x}=\int_{0}^{1}|f(\bm{\gamma}(t))|\|\bm{\gamma}^{\prime}(t)\|dt$. For a ReLU network (Eq.˜1), computing the error over the regions $\Omega_{\bm{\gamma}}=\Omega\cap S^{1}$ yields $e=2\pi\sum\limits_{k=0}^{r^{\prime}}\left((-1)^{z_{k}(t_{k})}\left[g_{r_{k}}(t)\right]_{t_{k}}^{s_{k}}+(-1)^{z_{k}(t_{k+1})}\left[g_{r_{k}}(t)\right]_{s_{k}}^{t_{k+1}}\right)$ for $r^{\prime}=|\Omega_{\bm{\gamma}}|$, where $t_{k}$ are the spline breakpoints pulled-back from $\mathbb{R}^{2}$ to [0,1]{[}0,1{]}, $r_{k}$ denotes which spline region the $k$-th segment $t_{k},t_{k+1}]$ falls into, $z_{k}(t):=\mathbf{1}_{\{\mathbf{A}_{r_{k},\cdot}\bm{\gamma}(t)+\mathbf{b}_{r_{k}}<0\}}$, $s_{k}\in[t_{k},t_{k+1}]$ is the value flip point of $z_{k}(t)$ and $g_{r_{k}}(t)=\mathbf{A}_{r_{k},1}\frac{\sin 2\pi t}{2\pi}-\mathbf{A}_{r_{k},2}\frac{\cos 2\pi t}{2\pi}+\mathbf{b}_{r_{k}}t$ (full derivations in Section˜C.2). Assuming the network considered attains optimal approximation error $e>0$, then it is clear that $t_{k+1}\to t_{k}\implies e\to 0$. The converse implication holds when the task is changed from classification to regressing the unit circle. In this case, reducing the approximation error by adapting the model’s weights (either through PEFT or training a larger model from scratch) will still result in an affine spline operator, for which $e>0$ and further reducing $e$ to 0 would require infinitely many neurons. This paper explores an orthogonal approach: by modulating the model’s activation functions, one can efficiently control its curvature and, in turn, that of its decision boundaries, thereby steering them toward optimality—without modifying the model’s weights. Fig.˜2 illustrates this phenomenon: methods such as LoRA implicitly tune the spline slopes and breakpoints, whereas modulating model nonlinearities changes the model’s underlying geometry.
3.1 The $\beta$-Vector-Quantization (VQ) inference framework
This section builds upon the max-affine spline formulation from Eq.˜2 to construct a model steering method operating on the model’s activation functions. By inspecting Eq.˜2, we observe that the mapping remains affine within each (implicitly defined) region where the pointwise maximum does not change. Specifically, for any input $\mathbf{x}$ where $\arg\max_{r=1\dots R}\big(\langle\mathbf{A}_{r,\cdot},\mathbf{x}\rangle+\mathbf{b}_{r}\big)$ remains constant, all such inputs belong to the same region, as they share the same affine mapping. The nonlinearity of the function arises when transitioning between these regions.
Smoothing the nonlinearity by smoothing the spline region assignment process. Instead of going from one affine mapping to another in an abrupt fashion (whenever crossing that hyperplane), one may consider a smoother transition. There are two common practices to achieve that goal.
@@ 3.1 The $\beta$-Vector-Quantization (VQ) inference framework
1. VQ Inference Step (region selection): Determine the affine transformation that maximizes the output, which can be viewed as a VQ process. The decision is encoded in a one-hot selection variable $\mathbf{t}\in\mathbb{R}^{R}$, where $R$ is the number of input region partitions of the max-affine spline function, and the $r^{*}$-th entry is set to 1, where:
@@
Using the computation step in Eq.˜4 and a ReLU activation function, switching from $\beta=1$ to $\beta=0.5$ is provably equivalent to replacing ReLU with the Sigmoid Linear Unit (SiLU). In the limit as $\beta\to 0$, the activation function becomes linear—thus making the entire input-output mapping of the network linear as well. [1]
Smoothing the nonlinearity by smoothing the max. As previously mentioned, there is an alternative way to smooth the max-affine spline mapping from Eq.˜2. Instead of relying on a soft region assignment, we can instead directly smooth the maximum function. It is already well known that smoothing the maximum operator leads to the log-sum-exp operator (i.e. Softplus). Hence, the mapping from Eq.˜2 now becomes
@@
where we parameterized the mapping so that its behavior is akin to Eq.˜5, a value of $\beta\to 1$ recovers the original affine spline activation, e.g., ReLU.
The crucial observation we make is that both parameterizations tend to shift the mean of the output of the unit either by a negative factor (for Eq.˜5) or a positive factor (for Eq.˜7). This means that in very deep models, varying $\beta$ with either parameterization produces a shift in the decision boundary or regression that cannot be recovered unless the parameters are trained once again, which we are trying to avoid. As a result, as detailed in Section˜3.2, our implementation combines the two parameterizations in a weighted manner to mitigate this bias, as Fig.˜S1 illustrates.
@@ 3.2 Implementation of CT
We begin by presenting the core activation that gives CT its expressive power—referred to as CT Unit (CTU). The activation is obtained by combining the two parameterizations discussed in Section˜3.1, in order to mitigate the mean shift introduced by each parameterization individually:
@@
It is worth noting that CTU naturally encompasses a broad family of functions. Particularly, it recovers SiLU ($c=1$), Softplus ($c=0$), as well as closely approximates GELU ($c=1$ and $\beta=0.64$).
Steering vs Trainable CT. We provide two implementations of CT differing in how CTU is applied. The first, denoted Steering CT (S-CT), replaces all ReLUs in the network with CTUs using a fixed $c=0.5$ and a shared $\beta\in[0,1]$. This version is highly parameter-efficient—introducing only a single hyperparameter—and does not require backpropagation, making it suitable as a steering method.
The second, referred to as Trainable CT (T-CT), also replaces all ReLUs with CTUs but assigns each output neuron its own trainable pair $(\beta,c)$, optimized via backpropagation. This version serves as a finetuning method: while it introduces additional parameters, the increase is modest compared to methods like LoRA and it yields competitive performance, as shown in Section˜4.2. Code for both implementations is provided in Appendix˜D.